Equivalent Method; 1. 4x-6y=6 ; 7x-3y=6 2. 3x-3y=8; 9x- 4y=3


1
Expert's answer
2022-03-16T04:05:59-0400

"1. \\
4x-6y=6\\
7x-3y=6\\
\left\{\begin{matrix}
 4x-6y=6\\
7x-3y=6
\end{matrix}\right.\\\left\{\begin{matrix}
 4x-6y=6\\
-14x+6y=-12
\end{matrix}\right.\\
-10x=-6\\
x=0.6\\
7\cdot0.6-3y=6\\
3y=-1.8\\
y=-0.6\\
(0.6,-0.6)"

"2.\\3x-3y=8\\
9x-4y=3\\
\left\{\begin{matrix}
 3x-3y=8\\
9x-4y=3
\end{matrix}\right.\\
\left\{\begin{matrix}
 -9x+9y=-24\\
9x-4y=3
\end{matrix}\right.\\
5y=-21\\
y=-\frac{21}{5}\\
3x-3\cdot(-\frac{21}{5})=8\\
3x=-\frac{23}{5}\\
x=-\frac{23}{15}\\
(-\frac{23}{15},-\frac{21}{5})"